1. How to calculate area of a triangle
1
Expert's answer
2021-09-14T00:06:18-0400

When we know two sides "a" and "b" of a triangle and the included angle "\alpha" to calculate the area "A" of a triangle we can use the following formula:

"A=\frac{1}{2}ab\sin \alpha."